Reading from left to right, the colored bands on a resistor are violet(7), yellow (4), brown(1), and gold(±5%). If the resistor measures 704 with an ohmmeter, it is


1
Expert's answer
2021-05-20T20:14:42-0400

Gives

Band Color code

Violet=7

Yellow =4

Brown =1

Gold=5%



Resistance measure

"R=74\times10^1+(74\times10^1)\times5" %

"R=74\times10^1\mp(74\times10^1)\times\frac{5} {100}"

"R=740\mp740\times\frac{5}{100}"

"R=740\mp37"

"R_{max}=777\Omega"

"R_{min}=703\Omega"